OpenStudy (anonymous):

Find the area A of a rectangle with length 8.9 ft and width 10.8 ft. Question 20 Find the area A of a rectangle with length 8.9 ft and width 10.8 ft.

hero (hero):

A = 8.9*10.8

OpenStudy (matt101):

A=l*w=8.9*10.8=96.12 ft^2

OpenStudy (anonymous):

Hi Hero, are you sure about this answer? I think that is A=l*w=8.9*10.8=96.12 ft^2

hero (hero):

I just gave him a medal

OpenStudy (anonymous):

That is means that matt10 is right...then